# Reinforcement Learning Pair Trading: A Dynamic Scaling Approach

### v1

@@ Abstract
Cryptocurrency is a cryptography-based digital asset with extremely volatile prices. Around \$70 billion worth of crypto-currency is traded daily on exchanges. Trading crypto-currency is difficult due to the inherent volatility of the crypto-market. In this work we want to test the hypothesis: “can techniques from artificial intelligence help with algorithmically trading crypto-currencies”. In order to address this question; we combine Reinforcement Learning (RL) with pair trading. Pair trading is an statistical arbitrage trading technique, which exploits the price difference between statistically correlated assets. We train reinforcement learners to determine when and how to trade pairs of crypto-currencies. We develop new reward shaping and observation/action spaces for reinforcement learning. We performed experiments with the developed reinforcement learner on pairs of BTC-GBP and BTC-EUR data separated by 1-minute intervals ($n=263,520$). The traditional non-RL pair trading technique achieved annualised profit of 8.33%, while the proposed RL based pair trading technique achieved annualised profits from 9.94% — 31.53%, depending upon the RL learner. Our results show that RL can significantly, outperform manual and traditional pair trading techniques when applied to volatile markets such as crypto-currencies.
@@
Pair Trading , Reinforcement Learning , Algorithmic Trading , Deep Learning , Cryptocurrency
journal: Expert Systems with Applications
@@ 1 Introduction
Arbitrage is a subdomain of financial trading that profits from price discrepancies in different markets [Dybvig & Ross, 1989]. Pair trading is one of the well-known arbitrage trading methods in financial markets. Arbitrageurs identify two highly correlated assets to form a pair. When a price discrepancy happens, they buy the underpriced asset and sell the overpriced correlated asset to profit from the mean reversion of the prices. Arbitrage requires practitioners to constantly analyse the market conditions at the fastest speed possible, as arbitrageurs must compete for transitory opportunities [Brogaard et al., 2014]. The faster the market analysis is, the more the chances of profiting from arbitrage. Therefore, we want to explore the process of utilising Artificial Intelligence (AI)to accelerate the process of pair trading.
@@
Reinforcement Learning (RL)is a captivating domain of AI. The idea of RL is to let the agent(s) learn to interact with an environment. The agent should learn from the environment’s responses to optimise its behaviour [Sutton & Barto, 2018]. If we view the financial market from the perspective of RL environment, actions in the financial market are investment decisions. To gain profits, arbitrageurs are incentivized to train agents that produce lucrative investment decisions, and RL facilitates agents’ learning process from the profit/loss of the market.
@@
The combination of RL and various financial trading techniques is still going through rapid evolution. There have been some work in RL infrastructural construction [Liu et al., 2021, 2022b, 2022a], and some experiments in profitable RL agent training [Meng & Khushi, 2019, Zhang et al., 2020, Pricope, 2021]. Trading actions in traditional pair trading follow static rules. In reality, the complexity of financial markets should allow more flexibility in the decision-making process. An experienced trader might analyse the market conditions to make informed decisions. However, it is not feasible to output efficient decisions at short, intermittent intervals 24/7. RL algorithms enable a fast-track decision-making process for analysing trading signals and generating trading actions.
@@
In designing a high-frequency trading system based on RL, several problems must be explored to ensure a fast decision-making process. The first problem is pair formation: identifying compatible instruments with historical correlations to form profitable pairs. The second problem concerns timing: instead of blindly following preset rules, the system needs flexibility in choosing investment timing for greater profit. The last challenge involves investment quantity: as investment opportunities vary in quality, experienced traders can select better opportunities with stronger profit potential. It is worth investigating whether RL agent is capable of achieving similar profitability by scrutinizing each investment opportunity.
@@
This paper investigates some questions centred around RL in pair trading. To overcome the fast decision-making requirement in a high-frequency trading environment, we constructed an environment that suits the RL agent to conduct pair trading and fine-tune reward shaping to encourage the agent to make profitable decisions. The contributions of this paper are: (i) proposal of a novel pair trading method that is adaptive to high volatility markets. (ii) utilisation of grid search technique to fine-tune profitable hyperparameters in pair trading. (iii) introduction of the RL component in pair trading for market analysis and decision-making. (iv) development of a novel RL model for making decisions about the quantity of investment.
@@
The structure of the paper is arranged as follows: the background and related work are introduced in Sections 2 and 3. The methodology is presented in Section 4. Experiments and results are included in Section 5. Following by discussion about the results and conclusion in Section 6.
@@ 2.1 Traditional Pair Trading
Classical pair trading consists of two distinct components known as legs. a leg represents one side of a trade in a multi-contract trading strategy. Under the definition of pair trading, “longing the first asset and shorting the second asset” is called long leg, and “shorting the first asset and longing the second asset” is called short leg. The two assets are always bought and sold in opposite directions in pair trading. Therefore, the overall pair trading strategy is considered to be market neutral, because the profits from the long position and the short are offset by the direction of the overall market. Gatev et al.’s 2006 work is the most cited traditional pair trading method. It follows the OODA (Observe, Orient, Decide and Act) Loop [Fadok et al., 1995]. Before entering the market, the first step is to choose the proper assets in a pair. Sum of Squared Deviation (SSD)is the measurement calculated from prices for assets $i$ and $j$. Through exhaustive searching in a formation period $T$, the assets with the smallest SSD are bound as a pair (Equation 1).
$SSD_{p_{i},p_{j}}=\sum_{t=1}^{T}(p_{i}-p_{j})^{2}$ (1)
  • 1.
@@
  • 2.
    Orient is exploring what could be done. Three possible actions for pair trading are long leg, short leg and close position as defined above.
  • 3.
@@
  • 4.
    Act once the decision is made. The long leg orders to buy asset $i$ and sell asset $j$; The short leg orders to sell asset $i$ and buy asset $j$. Closing a position means clearing all the active positions to hold cash only.
Figure 1: Price distance view of pair trading $p_{i}$ and $p_{j}$
A graphical visualisation of pair trading is presented in Figure 1. Figure 1 (a) is the market interactions according to the spread and thresholds. A position is opened whenever the spread deviates beyond the open threshold. The position closure happens when the spread reverts below the close threshold. Figure 1 (b) presents the corresponding actions with the degree of spread deviation. The spread deviations are classified into zones based on the Spread ($S$), Open-Threshold ($OT$) and Close-Threshold ($CT$) (Equation 2):
$\left\{\begin{array}[]{rl}\text{Short Zone:}&+OT<S<+\infty,\\ &\text{spread deviates beyond open threshold}\\ \text{Neutral Short Zone:}&+CT<S<+OT,\\ &\text{spread deviates between open and close threshold}\\ \text{Close Zone:}&-CT<S<+CT,\\ &\text{spread reverts between close thresholds}\\ \text{Neutral Long Zone:}&-OT<S<-CT,\\ &\text{spread deviates between open and close threshold}\\ \text{Long Zone:}&-\infty<S<-OT,\\ &\text{spread deviates below open threshold}\\ \end{array}\right.$ (2)
@@ 2.2 Reinforcement Learning
Reinforcement Learning (RL)is used for training an agent for maximising rewards while interacting with an environment [Sutton & Barto, 2018]. The environment for RL is required to be a Markov Decision Process (MDP[Bellman, 1957], which means it is modelled as a decision-making process with the following elements 〈State ($S$), Action ($A$), Transition ($P_{A}$), Reward ($R_{A}$)⟩. The goal is to train the agent to develop a policy ($\pi$) that fulfil and objective, e.g., maximising profits in a trade. At every trading interval $t$, according to the state $S$, that the agent observes, an action $A$ is chosen based on the policy $\pi$. The environment rewards/punishes the state transition of $S_{t}\rightarrow S_{t+1}$ with environment reward $r$. If we assume $\gamma$ to be the discount factor for the time-value discount of future reward, RL trains a policy $\pi$ that maximises the total discounted reward $G_{t}$ as shown in Equation 3:
$G_{t}=\sum_{i=0}^{\infty}\gamma^{i}r_{t+i}$ (3)
RL algorithms can broadly be classified based on three criteria: value/policy-based, on/off-policy and actor/critic-network [AlMahamid & Grolinger, 2021]. Value-based methods estimate state-action value functions for decision-making. Policy-based methods directly learn action selection policies. The on-policy method requires data generated by the current policy, and off-policy is capable of leveraging past experiences from potentially different policies. Moreover, actor-critic architectures, where an actor-network proposes actions and a critic network evaluates, have shown a better performance in facilitating policy improvement through this feedback loop. Most recent researches favour actor-critic architecture instead of actor-only or critic-only methods for better performance [Meng & Khushi, 2019, Zhang et al., 2020]. Therefore, only actor-critic algorithms are adopted in this study.
Based on the RL classification criteria, some representative algorithms have been selected for this study including Deep Q-Learning (DQN)[Mnih et al., 2013], Soft Actor Critic (SAC)[Haarnoja et al., 2019], Advantage Actor-Critic (A2C)[Sutton & Barto, 2018], Proximal Policy Optimization (PPO)[Schulman et al., 2017]. Diversified RL algorithms are experimented with to choose the most effective one in pair trading.
@@ 3.1 Reinforcement Learning in Algorithmic Trading
Reinforcement Learning in AlphaGo captured the world’s attention since 2016 by participating in a series of machine versus human competitions on board game GO [David Silver, Demis Hassabis, 2016]. Surprisingly, the research regarding RL in the financial market started long before that. Recurrent reinforcement learning studies were the mainstream works [Gold, 2003, Bertoluzzo & Corazza, 2007, Maringer & Ramtohul, 2012, Zhang & Maringer, 2016] in the early stage of financial trading. After the upsurge of AlphaGo, some significant advancements were brought to RL trading as well, Huang re-described Markov Decision Process (MDP)financial market as a game process to incorporate RL as Financial Trading as a Game [Huang, 2018]. Newer RL models such as Deep Q-Learning (DQN), Policy Gradients (PG)and Advantage Actor-Critic (A2C)have also been used recently by researchers [Meng & Khushi, 2019, Zhang et al., 2020, Pricope, 2021] for financial trading. A noteworthy research work is that of FinRL group in the infrastructures and ensemble learning mechanism [Liu et al., 2021, 2022b, 2022a].
@@ 3.2 Reinforcement Learning in Pair Trading
Reinforcement Learning, in combination with pair trading, is not an untapped domain. RL has ameliorated multifaceted aspects from the traditional method of pair trading brought up by Gatev et al. 2006. RL technique Ordering Points To Identify The Clustering Structure (OPTICS)contributed at the pair selection stage by leveraging a clustering algorithm to produce better pair choices [Sarmento & Horta, 2020]. Vergara & Kristjanpoller 2024 brought deep reinforcement learning into the Cryptocurrency world in an ensemble setting. reCurrent Reinforcement lEarning methoD for paIrs Trading (CREDIT)algorithm that takes consideration of both profitability and risks was engineered by Han et al. 2023. Reward Shaping is also an interesting area where some work has been done in RL trading [Lucarelli & Borrotti, 2019, Wang et al., 2021]. Kim & Kim’s work 2019 is the most recent RL pair trading method. Their focus is on utilising RL to find most trading opportunities. Instead of fixed thresholds, the RL agent in Kim & Kim’s work produces the thresholds for the coming trading period. Open, close and stop-loss thresholds determine the profits of pair trading.
Our work introduces a novel method to combine RL with pair trading. The work of Gatev et al. is not efficient enough for a high-frequency market. The state-of-the-art method of Kim & Kim has some deficiencies: (i) it requires the market’s volatility to be relatively stable. The RL agent may produce unsuitable thresholds if the market experiences increased volatility. (ii) It lacks flexibility in the investment amount. Opportunities with different qualities are programmed to be invested with the same amount of capital. Once the RL agent determines a threshold, the trading algorithm executes a trade at pre-determined thresholds. We leverage RL to make investment timing and quantity decisions. The adjustable investment amount is a novel feature of our RL pair trading. RL agent measures how well the investment opportunities are based on observations and invests a larger amount in more promising market conditions. Having another dimension on the investment side should further enhance profitability and reduce risks.
@@ 4 Methodology
In this section, we introduce the architecture of the methodology (Figure 2). The architecture includes five steps: (1) Pair Formation for selecting assets to form a tradeable pair (Section 4.1); (2)Spread Calculation utilising the moving-window technique to extract the spread in a limited retrospective time frame (Section 4.2); (3) Parameter Selection from historical dataset to decide the most suitable hyperparameters for pair trading (Section 4.2); (4) RL Trading by allowing RL to decide the trading timing and quantity in pair trading (Section 4.4); (5) Investment Action for taking the actions produced from RL trading into market execution.

### v2

# pair trading; reinforcement learning; algorithmic trading; deep learning; cryptocurrency

@@ Abstract
Cryptocurrency is a cryptography-based digital asset with extremely volatile prices. Around USD 70 billion worth of cryptocurrency is traded daily on exchanges. Trading cryptocurrency is difficult due to the inherent volatility of the crypto market. This study investigates whether Reinforcement Learning (RL) can enhance decision-making in cryptocurrency algorithmic trading compared to traditional methods. In order to address this question, we combined reinforcement learning with a statistical arbitrage trading technique, pair trading, which exploits the price difference between statistically correlated assets. We constructed RL environments and trained RL agents to determine when and how to trade pairs of cryptocurrencies. We developed new reward shaping and observation/action spaces for reinforcement learning. We performed experiments with the developed reinforcement learner on pairs of BTC-GBP and BTC-EUR data separated by 1 min intervals ($n$ = 263,520). The traditional non-RL pair trading technique achieved an annualized profit of 8.33%, while the proposed RL-based pair trading technique achieved annualized profits from 9.94% to 31.53%, depending upon the RL learner. Our results show that RL can significantly outperform manual and traditional pair trading techniques when applied to volatile markets such as cryptocurrencies.
@@
pair trading; reinforcement learning; algorithmic trading; deep learning; cryptocurrency
[…]
@@ 1 Introduction
Arbitrage is a subdomain of financial trading that profits from price discrepancies in different markets (Dybvig and Ross, 1989). Pair trading is one of the well-known arbitrage trading methods in financial markets. Arbitrageurs identify two highly correlated assets to form a pair. When a price discrepancy happens, they buy the underpriced asset and sell the overpriced correlated asset to profit from the mean reversion of the prices. With the rise of high-frequency trading, the ability to conduct fast and accurate analyses has become critical. Arbitrage requires practitioners to constantly analyze the market conditions at the fastest speed possible, as arbitrageurs must compete for transitory opportunities (Brogaard et al., 2014). Therefore, we explore how Artificial Intelligence (AI)can enhance the process of pair trading, focusing on the speed and adaptability of decision-making.
@@
Reinforcement Learning (RL)is a captivating domain of AI. The idea of RL is to let the agent(s) learn to interact with an environment. The agent should learn from the environment’s responses to optimize its behavior (Sutton and Barto, 2018). If we view the financial market from the perspective of the RL environment, actions in the financial market are investment decisions. By allowing agents to adapt dynamically to market conditions, RL has the potential to overcome the limitations of static, rule-based strategies in volatile and complex financial environments. For gaining profits, arbitrageurs are incentivized to train agents to produce lucrative investment decisions, and RL facilitates agents’ learning process from the profit/loss of the market.
@@
The combination of RL and various financial trading techniques is still evolving rapidly. There has been some work in RL infrastructural construction (Liu et al. 2021, 2022a, 2022b) and some experiments in profitable RL agent training (Meng and Khushi, 2019; Zhang et al., 2020; Pricope, 2021). Trading actions in traditional pair trading follow static rules. In reality, the complexity of financial markets should allow more flexibility in the decision-making process. An experienced trader might analyze market conditions to make informed decisions. However, it is not feasible to output efficient decisions at short, intermittent intervals 24/7. RL algorithms enable a fast-track decision-making process for analyzing trading signals and generating trading actions.
@@
Designing a high-frequency trading system based on RL requires addressing critical challenges. The first challenge is how to construct an RL environment that accommodates RL agents for arbitrage. The second challenge involves identifying compatible instruments with historical correlations to form profitable pairs. The third challenge concerns timing. Instead of blindly following preset rules, the system requires flexibility in choosing investment timings to achieve greater profitability. The final challenge involves investment quantity. Since investment opportunities vary in quality, a critical consideration is whether RL agents can replicate decision-making capabilities comparable to the scrutiny applied by experienced traders.
@@
This paper investigates key questions centered around the application of Reinforcement Learning (RL) in pair trading. To address the fast decision-making requirements in a high-frequency trading environment, we constructed a tailored RL environment for pair trading and fine-tuned reward shaping to encourage the agent to make profitable decisions. The contributions of this work are as follows: (1) the construction of an RL environment specifically designed for quantity-varying pair trading; (2) the proposal of a novel pair trading method that incorporates adaptive investment quantities to capture opportunities in highly volatile markets; (3) the use of a grid search technique to fine-tune hyperparameters for enhanced profitability; (4) the introduction of an RL component for market analysis and decision-making in pair trading, along with a novel RL model optimized for investment quantity decisions.
@@
The structure of the paper is arranged as follows: the background and related work are introduced in Sections 2 and 3. The methodology is presented in Section 4. Experiments and results are included in Section 5. A discussion of the results and conclusions is provided in Section 6.
@@ 2.1 Traditional Pair Trading
Classical pair trading consists of two distinct components known as legs . A leg represents one side of a trade in a multi-contract trading strategy. Under the definition of pair trading, “longing the first asset and shorting the second asset” is called a long leg , and “shorting the first asset and longing the second asset” is called a short leg . The two assets are always bought and sold in opposite directions in pair trading. Therefore, the overall pair trading strategy is considered to be market neutral because the profits from the long position and the short are offset by the direction of the overall market. Gatev et al.’s (2006) work is the most cited traditional pair trading method. It follows the OODA (observe, orient, decide, and act) loop (Fadok et al., 1995). Before entering the market, the first step is to choose the proper assets in a pair. The Sum of Squared Deviation (SSD)is the measurement calculated from the prices of assets $i$ and $j$. Through exhaustive searching in a formation period $T$, the assets with the smallest SSD are bound as a pair Equation (1).
$SSD_{p_{i},p_{j}}=\sum_{t=1}^{T}(p_{i}-p_{j})^{2}.$ (1)
@@
  • Orient is the process exploring what could be done. Three possible actions for pair trading are long leg, short leg, and close position, as defined above.
@@
  • Act once the decision is made. The long leg orders us to buy asset $i$ and sell asset $j$. The short leg orders us to sell asset $i$ and buy asset $j$. Closing a position means clearing all the active positions to hold cash only.
A graphical visualization of pair trading is presented in Figure 1. Figure 1a shows the market interactions according to the Spread (S) and thresholds. A position is opened whenever the spread deviates beyond the open threshold. The position closure happens when the spread reverts below the close threshold. Figure 1b, which shares the same time axis with (a), is a stretched view of (a). It presents the corresponding actions with the crossing of Spread (S) and zones. The spread deviations are classified into zones based on the Spread ($S$), Open-Threshold ($OT$), and Close-Threshold ($CT$) Equation (2):
Figure 1: Stretched pair trading view of price distance between $p_{i}$ and $p_{j}$. Figure (b), which shares the same time axis with (a), is a stretched view of (a). It presents the corresponding same actions with the crossing of Spread (S) and zones in two different views.
$\left\{\begin{array}[]{rl}\text{Short Zone:}&+OT<S<+\infty,\\ &\text{spread deviates beyond open threshold}\\ \text{Neutral Short Zone:}&+CT<S<+OT,\\ &\text{spread deviates between open and close threshold}\\ \text{Close Zone:}&-CT<S<+CT,\\ &\text{spread reverts between close thresholds}\\ \text{Neutral Long Zone:}&-OT<S<-CT,\\ &\text{spread deviates between open and close threshold}\\ \text{Long Zone:}&-\infty<S<-OT,\\ &\text{spread deviates below open threshold}\\ \end{array}.\right.$ (2)
@@ 2.2 Reinforcement Learning
Reinforcement Learning (RL)is used to train an agent to maximize rewards while interacting with an environment (Sutton and Barto, 2018). The environment for RL is required to be a Markov Decision Process (MDP(Bellman, 1957), which means it is modeled as a decision-making process with the following elements〈State ($S$), Action ($A$), Transition ($P_{A}$), Reward ($R_{A}$)⟩. The goal is to train the agent to develop a policy ($\pi$) that fulfills an objective, e.g., maximizing profits in a trade. At every trading interval $t$, according to state $S$, that the agent observes, action $A$ is chosen based on policy $\pi$. The environment rewards/punishes the state transition of $S_{t}\rightarrow S_{t+1}$ with environment reward $r$. If we assume $\gamma$ to be the discount factor for the time-value discount of future reward, RL trains a policy $\pi$ that maximizes the total discounted reward $G_{t}$, as shown in Equation (3):
$G_{t}=\sum_{i=0}^{\infty}\gamma^{i}r_{t+i}.$ (3)
RL algorithms can broadly be classified based on three criteria: value/policy-based, on/off policy, and actor/critic network (AlMahamid and Grolinger, 2021). Value-based methods estimate state-action value functions for decision-making. Policy-based methods directly learn action selection policies. The on-policy method requires data generated by the current policy, and the off-policy method is capable of leveraging past experiences from potentially different policies. Moreover, actor–critic architectures, where an actor–network proposes actions and a critic network evaluates, have shown a better performance in facilitating policy improvement through this feedback loop. Most recent research favors an actor–critic architecture instead of actor-only or critic-only methods for better performance (Meng and Khushi, 2019; Zhang et al., 2020). Therefore, only actor–critic algorithms are adopted in this study.
Based on the RL classification criteria, some representative algorithms have been selected for this study including Deep Q-Learning (DQN)(Mnih et al., 2013), Soft Actor Critic (SAC)(Haarnoja et al., 2019), Advantage Actor-Critic (A2C)(Sutton and Barto, 2018), Proximal Policy Optimization (PPO)(Schulman et al., 2017). Diversified RL algorithms are experimented with to choose the most effective one in pair trading.
@@ 3.1 Reinforcement Learning in Algorithmic Trading
Reinforcement learning in AlphaGo captured the world’s attention in 2016 by participating in a series of machine versus human competitions on the board game GO (Silver and Hassabis, 2016). Surprisingly, the research regarding RL in the financial market started long before that. Recurrent reinforcement learning studies were the mainstream works (Gold, 2003; Bertoluzzo and Corazza, 2007; Maringer and Ramtohul, 2012; Zhang and Maringer, 2016) in the early stage of financial trading. After the upsurge of AlphaGo, some significant advancements were brought to RL trading as well; Huang (2018) re-described the Markov Decision Process (MDP)financial market as a game process to incorporate RL as a financial trading game (Huang, 2018). Pricope (2021) proposed deep RL agents to develop profitable high-frequency trading strategies with sequential model-based optimization tuning the hyperparameters. With the recent development, newer RL models such as Deep Q-Learning (DQN), Policy Gradients (PG), and Advantage Actor-Critic (A2C)have also been introduced by researchers (Meng and Khushi, 2019; Zhang et al., 2020; Mohammadshafie et al., 2024) for financial trading. A noteworthy research work is that of the FinRL group in the infrastructures and ensemble learning mechanism (Liu et al. 2021, 2022a, 2022b).
@@ 3.2 Reinforcement Learning in Pair Trading
Reinforcement Learning, in combination with pair trading, is not an untapped domain. RL has ameliorated multifaceted aspects of the traditional method of pair trading brought up by Gatev et al. (2006). The RL technique Ordering Points To Identify The Clustering Structure contributed to the pair selection stage by leveraging a clustering algorithm to produce better pair choices (Sarmento and Horta, 2020). Vergara and Kristjanpoller (2024) traded deep reinforcement learning in the cryptocurrency market with ensemble practice, combining classical pair trading with the RL framework. reCurrent Reinforcement lEarning methoD for paIrs Trading (CREDIT), an algorithm that takes into consideration both profitability and risks, was engineered by Han et al. (2023). Reward shaping is also an interesting area where some work has been done in RL trading (Lucarelli and Borrotti, 2019; Wang et al., 2021). Kim and Kim’s work in (2019) is the most recent RL pair trading method. Their focus is on utilizing RL to find the most trading opportunities. Instead of fixed thresholds, the RL agent in Kim and Kim’s work produces thresholds for the upcoming trading period. Open, close, and stop-loss thresholds determine the profits of pair trading.
Our work introduces a novel method to combine RL with pair trading. The work of Gatev et al. (2006) is not efficient enough for a high-frequency market. The state-of-the-art method of Kim and Kim (2019) has some deficiencies: (i) it requires the market’s volatility to be relatively stable. The RL agent may produce unsuitable thresholds if the market experiences increased volatility. (ii) It lacks flexibility in the investment amount. Opportunities with different qualities are programmed to be invested with the same amount of capital. Once the RL agent determines a threshold, the trading algorithm executes a trade at pre-determined thresholds. We leverage RL to make investment timing and quantity decisions. The adjustable investment amount is a novel feature of our RL pair trading. An RL agent measures how well the investment opportunities are based on observations and invests a larger amount in more promising market conditions. Having another dimension on the investment side should further enhance profitability and reduce risks.
@@ 4 Methodology
In this section, we introduce the architecture of the methodology (Figure 2). The architecture includes five steps: (1) pair formation for selecting assets to form a tradeable pair (Section 4.1); (2) spread calculation utilizing the moving-window technique to extract the spread in a limited retrospective time frame (Section 4.2); (3) parameter selection from an historical dataset to decide the most suitable hyperparameters for pair trading (Section 4.2); (4) RL trading by allowing RL to decide the trading timing and quantity in pair trading (Section 4.4); (5) investment action for taking the actions produced from RL trading into market execution.